OpenStudy (anonymous):

HELP will fan and medal I need help understanding how I would write (and solve) the answer to this question: Using complete sentences, write an indirect proof proving that if x = 30, then 3 over 4x + 5 ≠ 20.

OpenStudy (anonymous):

Wait are you asking if 20 is the right answer?

OpenStudy (anonymous):

x = 30 , the 3 / 4 x + 5 (insert not equal sign) 20

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so basically i just have to write how to solve it? or?

OpenStudy (anonymous):

@MakaylaTracy @mathmath333 help please. if you're able

OpenStudy (mathmath333):

for indirect proof lets assume that \(\large\tt \begin{align} \color{black}{4x + 5 = 20 }\end{align}\) then \(\large\tt \begin{align} \color{black}{4x + 5 = 20\\~\\~\\ 4x=20-5=15\\~\\~\\ \bf x=\dfrac{15}{4}\approx 3.7 }\end{align}\) but this contradicts thr given data that \(\large\tt \begin{align} \color{black}{\bf x=30 }\end{align}\) hence \(\Large\tt \begin{align} \color{black}{4x + 5 \neq 20. }\end{align}\)

OpenStudy (anonymous):

Ohh. okay. So because 4x+5= 30 the statement given was correct?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

okay wait. i just reviewed it again. so because it does not equal 30 the statement was correct

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{4x+5\\ =4\times 30+5\\ =125\\ }\end{align}\) so \(\huge\tt \begin{align} \color{black}{4x+5=125 }\end{align}\)

OpenStudy (anonymous):

Okay. Would I follow this same format for the equation 3/4x + 5 ≠ 20 Thank you so much by the way, this is all starting to make sense.

OpenStudy (anonymous):

\[3/ 4x + 5 ≠ 20\] ***

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

rewrite the full question

OpenStudy (anonymous):

write an indirect proof proving that if x = 30 \[3/4x +5\neq20\] ^ full question

OpenStudy (mathmath333):

ok apply the same method as i stated before